Ruth is twice as old as Pat. In 4 years Ruth will be three times as old as Pat was 3 years ago. How old is each now?
Answer by mananth(16946)

Ruth is twice as old as Pat. In 4 years Ruth will be three times as old as Pat was 3 years ago. How old is each now?
..
let Pat's age be x years
Ruth's age will be 2x
...
after 4 years ruth will be 2x+4
...
3 years ago pat was x-3 years
...
2x+4=3(x-3)
2x+4=3x-9
2x-3x=-9-4
-x=-13
x= 13 years Pat's age.
Ruth's age = 2x = 26 years
...
m.ananth@hotmail.ca